跳到论文正文
W Papers独立论文阅读器 arXiv:2105.09831 论文正文 计算阅读时间 · 0%
引用这篇论文根据当前论文元数据生成,正式投稿前建议核对。

    
  
License: arXiv.org perpetual non-exclusive license
arXiv:2105.09831v1 [cs.IT] 20 May 2021

Uncoded Binary Signaling through Modulo AWGN Channel

Gizem Tabak and Andrew Singer Affiliation: Coordinated Science Laboratory, Department of Electrical and Computer Engineering Affiliation: University of Illinois at Urbana–Champaign, Urbana, IL 61801 USA
Abstract

Modulo-wrapping receivers have attracted interest in several areas of digital communications, including precoding and lattice coding. The asymptotic capacity and error performance of the modulo AWGN channel have been well established. However, due to underlying assumptions of the asymptotic analyses, these findings might not always be realistic in physical world applications, which are often dimension- or delay-limited. In this work, the optimum ways to achieve minimum probability of error for binary signaling through a scalar modulo AWGN channel is examined under different scenarios where the receiver has access to full or partial information. In case of partial information at the receiver, an iterative estimation rule is proposed to reduce the error rate, and the performance of different estimators are demonstrated in simulated experiments.

††footnotetext: Appeared at the 54th IEEE Asilomar Conference on Signals, Systems & Computers.

I Introduction

Modulo wrapping is encountered in various fields of digital communications. In precoding literature, modulo operation is used to cancel out known intersymbol interference at the transmitter [1]. In information theory literature, nested lattice codes [2] and lattice Gaussian codes [3] are proven to achieve additive white Gaussian noise (AWGN) channel capacity asymptotically under certain conditions. Both works, and many more in lattice coding literature, utilize modulo operation on lattices to encode, quantize or decode data. Recently, modulo wrapping receiver front-ends are used in analog-to-digital conversion to decrease distortion or to reduce number of quantization bits [4].

The asymptotic capacity and the error performance of the modulo channel have been well studied under the existing coding literature [2, 5, 6]. The theoretical results, which are as impressive as achieving AWGN channel capacity asymptotically, are built on assumptions that are valid in high dimensions or long data block lengths. These assumptions might not always find correspondance in real physical applications which are often delay-limited and mostly built with hardware that operates in low dimensions. Moreover, in [4], it is demonstrated that already-existing voltage-controlled oscillator (VCO) ADCs can be utilized as modulo receiver front-ends. Hence, the realization of modulo channels in physical, practical applications is feasible. Therefore, the non-asymptotical communication performance of modulo channels in low dimensions is of increased interest. In this work, the aim is to develop an optimum way (in terms of decoding error) for binary signaling through a scalar modΔ\md AWGN channel with a power constraint, analyze the performance with different decision rules based on available information at the receiver, and propose an estimator to improve performance when the information at the receiver is lacking.

In Section II, the transmission and decoding scheme over modΔ\md AWGN channel will be introduced. In Section III, the optimum transmission scheme and the performance of the communication will be derived when the prior information is available at the receiver. In Section IV, first, the optimum transmission scheme and the decoding performance with the blind decision rule will be derived when the prior information is not available at the receiver. Then, an iterative estimation rule will be proposed to improve the performance. The error rates of different estimators will be demonstrated on simulated experiments in Section V, and the findings will be concluded in VI.

II Memoryless Mod-Δ\Delta Channel and Decoding

In this paper, the following communication system will be considered: NN uncoded, independent and identically distributed (i.i.d.) sequence of data bits a0N−1=(a0,…,an,…,aN−1)∈ℤ2N\textbf{a}_{0}^{N-1}=(a_{0},\dots,a_{n},\dots,a_{N-1})\in\mathbb{Z}_{2}^{N} are mapped into symbols x0M−1=(x0,…,xn,…,xM−1)∈𝒮M\textbf{x}_{0}^{M-1}=(x_{0},\dots,x_{n},\dots,x_{M-1})\in\mathcal{S}^{M} to be transmitted through a modulo AWGN channel. The modΔ\md AWGN channel is defined with the input-output relation

y~n=yn​modΔ=(xn+wn)​modΔ\tilde{y}_{n}=y_{n}\md=(x_{n}+w_{n})\md

where wnw_{n} are i.i.d. ∼fW​(w)=𝒩⁡(0,σ2)\sim f_{W}(w)=\mathcal{N}(0,\sigma^{2}) noise samples. The modulo operation modΔ:ℝ→[−Δ2,Δ2)\mod\Delta:\mathbb{R}\rightarrow\left[-\frac{\Delta}{2},\frac{\Delta}{2}\right) is defined as

xmodΔ=a⇔x=a+kΔ,k∈ℤ and a∈[−Δ2,Δ2).x\md=a\Leftrightarrow x=a+k\Delta,k\in\mathbb{Z}\text{ and }a\in\left[-\tfrac{\Delta}{2},\tfrac{\Delta}{2}\right).

In the binary signaling problem, the symbols are given by

xn={h0​, ​an=0h1​, ​an=1x_{n}=\begin{cases}h_{0}\text{, }a_{n}=0\\ h_{1}\text{, }a_{n}=1\end{cases} (1)

where 𝒮={h0,h1}\mathcal{S}=\{h_{0},h_{1}\} and the power is assumed P≥Δ2/2P\geq{\Delta^{2}}/{2}.

The objective is to find the optimum mapping in (1) that minimizes the probability of error with minimum power PP when recovering the sequence of data bits a0N−1\textbf{a}_{0}^{N-1} given the modulo-received samples y~0N−1\tilde{\textbf{y}}_{0}^{N-1}. More formally, minimizing the sequence error rate (SER) can be formulated as

a^0N−1=\displaystyle\hat{\textbf{a}}_{0}^{N-1}= argmaxa∈ℤ2NP​(a0N−1=a|y~0N−1)\displaystyle\argmax_{\textbf{a}\in\mathbb{Z}_{2}^{N}}P(\textbf{a}_{0}^{N-1}=\textbf{a}|\tilde{\textbf{y}}_{0}^{N-1})
=\displaystyle= argmaxa∈ℤ2NP​(y~|x=ha)​P​(x=ha)\displaystyle\argmax_{\textbf{a}\in\mathbb{Z}_{2}^{N}}P(\tilde{\textbf{y}}|\textbf{x}=\textbf{h}_{\textbf{a}})P(\textbf{x}=\textbf{h}_{\textbf{a}}) (2)

where the nn-th element of ha∈ℝN\textbf{h}_{\textbf{a}}\in\mathbb{R}^{N} is hanh_{a_{n}}. The subscripts and superscripts are omitted for the sake of simplicity.

If the priors P⁡(x=ha)P(\textbf{x}=\textbf{h}_{\textbf{a}}) are known at the receiver, then (II) corresponds to the maximum a posteriori (MAP) estimation. If the priors are not known, then the maximum likelihood (ML) estimation

a^0N−1=argmaxa∈ℤ2NP​(y~|x=ha)\displaystyle\hat{\textbf{a}}_{0}^{N-1}=\argmax_{\textbf{a}\in\mathbb{Z}_{2}^{N}}P(\tilde{\textbf{y}}|\textbf{x}=\textbf{h}_{\textbf{a}})

is used.

III Modulation and Decoding with Prior Information

If the prior information is available at the receiver, than the MAP estimator in (II) would be used for decoding to minimize the probability of error on the modΔ\md AWGN channel. Knowing the decision rule, the transmitter can optimize the symbols to be sent through the channel to minimize the probability of error and also the transmit power.

Define fWΔ:ℝ→ℝ+f_{W}^{\Delta}\colon\mathbb{R}\to\mathbb{R^{+}}, such that x↦fWΔ​(x)=∑k=−∞∞fW​(x+k​Δ)x\mapsto f_{W}^{\Delta}(x)=\sum_{k=-\infty}^{\infty}f_{W}(x+k\Delta) where W∼𝒩⁡(0,σ2)W\sim\mathcal{N}(0,\sigma^{2}). Then, the MAP estimation (II) yields

a^n\displaystyle\hat{a}_{n} =argmaxa∈ℤ2P​(y~n|an=a)​P​(an=a)\displaystyle=\argmax_{{a}\in\mathbb{Z}_{2}}P(\tilde{y}_{n}|a_{n}=a)P(a_{n}=a)
=argmaxa∈ℤ2∑k=−∞∞fW(y~n−ha+kΔ)P(an=a)\displaystyle=\argmax_{{a}\in\mathbb{Z}_{2}}\!\sum_{k=-\infty}^{\infty}f_{W}(\tilde{y}_{n}-h_{a}+k\Delta)P(a_{n}=a)
=argmaxa∈ℤ2fWΔ​(y~n−ha)​P​(an=a)\displaystyle=\argmax_{{a}\in\mathbb{Z}_{2}}f_{W}^{\Delta}(\tilde{y}_{n}\!-\!h_{a})P(a_{n}=a)

for n=0,…,N−1n=0,\dots,N-1, and the corresponding MAP decision rule, denoted as δπ0​(y~n)\delta_{\pi_{0}}(\tilde{y}_{n}), becomes

δπ0​(y~n)\displaystyle\delta_{\pi_{0}}(\tilde{y}_{n}) ={0,π0​fWΔ​(y~n−h0)≥(−π0)​fWΔ​(y~n−h1)1,π0​fWΔ​(y~n−h0)<(−π0)​fWΔ​(y~n−h1)\displaystyle\!=\!\begin{cases}0,\pi_{0}f_{W}^{\Delta}(\tilde{y}_{n}\!-\!h_{0})\!\geq\!(1\!-\!\pi_{0})f_{W}^{\Delta}(\tilde{y}_{n}\!-\!h_{1})\\ 1,\pi_{0}f_{W}^{\Delta}(\tilde{y}_{n}\!-\!h_{0})\!<\!(1\!-\!\pi_{0})f_{W}^{\Delta}(\tilde{y}_{n}\!-\!h_{1})\end{cases} (3)

where π0=P⁡(an=0)\pi_{0}=P(a_{n}=0).

The likelihood ratio for i.i.d. random bits an{a_{n}} is defined as

Lπ0​(an|y~n)=P⁡(an=0|y~n)P⁡(an=1|y~n)=P⁡(y~n|an=0)P⁡(y~n|an=1)​π01−π0L_{\pi_{0}}(a_{n}|\tilde{y}_{n})=\dfrac{P(a_{n}=0|\tilde{y}_{n})}{P(a_{n}=1|\tilde{y}_{n})}=\dfrac{P(\tilde{y}_{n}|a_{n}=0)}{P(\tilde{y}_{n}|a_{n}=1)}\dfrac{\pi_{0}}{1-\pi_{0}} (4)

and the MAP decision rule in (3) can be written as

a^n=δπ0​(y~n)\displaystyle\hat{a}_{n}=\delta_{\pi_{0}}(\tilde{y}_{n}) ={0,Lπ0​(an|y~n)≥11,Lπ0​(an|y~n)<1\displaystyle=\begin{cases}0,\>\>L_{\pi_{0}}(a_{n}|\tilde{y}_{n})\geq 1\\ 1,\>\>L_{\pi_{0}}(a_{n}|\tilde{y}_{n})<1\end{cases} (5)

and the probability of bit error is

Pe​(δπ0)=\displaystyle P_{e}(\delta_{\pi_{0}})= P⁡(a^n=1|an=0)​P​(an=0)\displaystyle P(\hat{{a}}_{n}=1|a_{n}=0)P(a_{n}=0)
+P⁡(a^n=0|an=1)​P​(an=1)\displaystyle+P(\hat{{a}}_{n}=0|a_{n}=1)P(a_{n}=1)

III-A Uniform Priors

Fig. 1: Likelihood functions and MAP decision regions. Probability of error with MAP decision rule, Pe​(δπ0)P_{e}(\delta_{\pi_{0}}), corresponds to the shaded area under the curves

Now, assume that the data bits are not only i.i.d., but also uniformly distributed (UID). In that case, the likelihood ratio in (4) becomes

L⁡(an|y~n)=P⁡(y~n|an=0)P⁡(y~n|an=1)L(a_{n}|\tilde{y}_{n})=\dfrac{P(\tilde{y}_{n}|a_{n}=0)}{P(\tilde{y}_{n}|a_{n}=1)}

and it results in the decision rule δ⁡(y~n)\delta(\tilde{y}_{n})

a^n=δ⁡(y~n)={0,L⁡(an|y~n)≥11,L⁡(an|y~n)<1\hat{a}_{n}=\delta(\tilde{y}_{n})=\begin{cases}0,\>\>L(a_{n}|\tilde{y}_{n})\geq 1\\ 1,\>\>L(a_{n}|\tilde{y}_{n})<1\end{cases}

Note that fWΔ​(y~n−ha)=fWΔ​(y~n−ha+k​Δ)​∀k∈ℤf_{W}^{\Delta}(\tilde{y}_{n}-h_{a})=f_{W}^{\Delta}(\tilde{y}_{n}-h_{a}+k\Delta)\>\forall k\in\mathbb{Z}, and {ha=h+k​Δ|k∈ℤ}\{h_{a}=h+k\Delta|\>k\in\mathbb{Z}\} yield the same decision and hence the same probability of error. Therefore, the transmitter can restrict the symbols to the interval [−Δ2,Δ2)\itv to reduce the average transmit power PP. Furthermore, the transmitter can set the values for hanh_{a_{n}} appropriately to minimize the probability of error, 1−P⁡(a^=a)=1−∏P⁡(a^n=an)1-P(\hat{\textbf{a}}=\textbf{a})=1-\prod P(\hat{a}_{n}=a_{n}), for given Δ\Delta and σ\sigma. This claim is formalized in Theorem III.1 and the proof is given in Appendix C.

Lemma 1.

If z∈ℝz\in\mathbb{R} and ℋ={c∈[−Δ2,Δ2)|fWΔ​(c)=z}\mathcal{H}=\left\{c\in\itv\middle|f_{W}^{\Delta}(c)=z\right\}, then |ℋ|≤2\left|\mathcal{H}\right|\leq 2.

Lemma 2.

Let aa be UID over ℤ2\mathbb{Z}_{2}, h0≠h1h_{0}\neq h_{1}, and define 𝒞={c∈[−Δ2,Δ2)|L⁡(a=0|c)=L⁡(a=1|c)}\mathcal{C}=\{c\in\itv\>\big|\>L(a=0|c)=L(a=1|c)\}. Then, 𝒞={h0+h12,h0+h12+sgn⁡(h0+h12)​Δ2}\mathcal{C}=\left\{\tfrac{h_{0}+h_{1}}{2},\tfrac{h_{0}+h_{1}}{2}+\sgn\left({\frac{h_{0}+h_{1}}{2}}\right)\tfrac{\Delta}{2}\right\} where sgn\sgn denotes the modified sign function so that sgn⁡(0)=−1\sgn(0)=-1.

Theorem III.1.

Let ana_{n} be UID over ℤ2\mathbb{Z}_{2} and xnx_{n} the symbols representing ana_{n} as defined in (1). Then, h0=−Δ4h_{0}=-\tfrac{\Delta}{4} and h1=Δ4h_{1}=\tfrac{\Delta}{4} yields the minimum probability of error, Pe∗P_{e}^{*}, while minimizing the average power P=𝔼⁡[xn2]P=\mathbb{E}[x_{n}^{2}] for Pe∗P_{e}^{*}.

If the symbol mapping in Theorem III.1 is used at the transmitter, then the probability of bit error becomes

Pe\displaystyle P_{e} =Φ⁡(Δ4​σ)−Φ⁡(3​Δ4​σ)\displaystyle=\Phi\left(\frac{\Delta}{4\sigma}\right)-\Phi\left(\frac{3\Delta}{4\sigma}\right)

where Φ⁡(xσ)=∑k=−∞∞Q⁡(x+k​Δσ)\Phi\left(\frac{x}{\sigma}\right)=\sum_{k=-\infty}^{\infty}Q\left(\frac{x+k\Delta}{\sigma}\right). Note that the probability of error with optimum transmit symbols depends on the ratio Δ/σ\Delta/\sigma.

III-B Nonuniform Priors

If the data bits are not UID but the prior π0\pi_{0} is known at the receiver, then the decision rule in (5) will be used to decode the received symbols. In Fig. 1, the likelihood functions, decision regions (horizontal bars) and the probability of error (shaded area) is demonstrated for Δ/σ=5\Delta/\sigma=5. To derive the symbol mapping that will yield the minimum probability of error with minimum average power when the priors are available at the receiver, first, define

P1\displaystyle P_{1} =π0​(Φ⁡(c1−h0)−Φ⁡(c2−h0))\displaystyle=\pi_{0}\left(\Phi(c_{1}-h_{0})-\Phi(c_{2}-h_{0})\right) (6)
+(1−π0)​(Φ⁡(c2−h1)−Φ⁡(c1−h1+Δ))\displaystyle+(1-\pi_{0})\left(\Phi(c_{2}-h_{1})-\Phi(c_{1}-h_{1}+\Delta)\right) (7)
P2\displaystyle P_{2} =π0​(Φ⁡(c2−h0)−Φ⁡(c1−h0+Δ))\displaystyle=\pi_{0}\left(\Phi(c_{2}-h_{0})-\Phi(c_{1}-h_{0}+\Delta)\right)
+(1−π0)​(Φ⁡(c1−h1)−Φ⁡(c2−h1))\displaystyle+(1-\pi_{0})\left(\Phi(c_{1}-h_{1})-\Phi(c_{2}-h_{1})\right)

where 𝒞={c1,c2}\mathcal{C}=\{c_{1},c_{2}\} is defined as in Lemma 1. Note that (6) corresponds to the red shaded region and (7) corresponds to the blue shaded region in Fig. 1. The, the probability of error can be written as

Pe​(δπ0)={P1,π0​fWΔ​(−Δ2−h0)>(1−π0)​fWΔ​(−Δ2−h1)P2,π0​fWΔ​(−Δ2−h0)≤(1−π0)​fWΔ​(−Δ2−h1)P_{e}(\delta_{\pi_{0}})\!=\!\begin{cases}P_{1},&\!\!\pi_{0}f_{W}^{\Delta}(-\frac{\Delta}{2}-h_{0})>(1-\pi_{0})f_{W}^{\Delta}(-\frac{\Delta}{2}-h_{1})\\ P_{2},&\!\!\pi_{0}f_{W}^{\Delta}(-\frac{\Delta}{2}-h_{0})\leq(1-\pi_{0})f_{W}^{\Delta}(-\frac{\Delta}{2}-h_{1})\end{cases}
Theorem III.2.

Let ana_{n} be i.i.d. over ℤ2\mathbb{Z}_{2} and xnx_{n} the symbols representing ana_{n} as defined in (1). Then, h0=−Δ4h_{0}=-\tfrac{\Delta}{4} and h1=Δ4h_{1}=\tfrac{\Delta}{4} yields the minimum probability of error with decision rule (3) while minimizing the average power P=𝔼⁡[xn2]P=\mathbb{E}[x_{n}^{2}].

Fig. 2: Probability of error with MAP decision rule Pe​(δπ0)P_{e}(\delta_{\pi_{0}}) with different priors π0\pi_{0} for different Δ/σ\Delta/\sigma ratios

Then, the corresponding probability of error becomes

Pe​(δπ0)=\displaystyle P_{e}(\delta_{\pi_{0}})= (8)
{1−π0,π0>11+τ(1−π0)​(Φ⁡(lπ0σ)−Φ⁡(Δ−lπ0σ))+π0​(Φ⁡(Δ/2−lπ0σ)−Φ⁡(Δ/2+lπ0σ)),11+τ≥π0>0.5(1−π0)​(Φ⁡(Δ/2−lπ0σ)−Φ⁡(Δ/2+lπ0σ))+π0​(Φ⁡(lπ0σ)−Φ⁡(Δ−lπ0σ)),τ1+τ≤π0<0.5π0,π0<τ1+τ\displaystyle\begin{cases}\hfil 1-\pi_{0}\!\!\!\!\!\!\!\!&\!\!\!\!\!,\pi_{0}>\frac{1}{1+\tau}\\ (1-\pi_{0})\left(\Phi\left(\frac{l_{\pi_{0}}}{\sigma}\right)-\Phi\left(\frac{\Delta-l_{\pi_{0}}}{\sigma}\right)\right)\\ +\pi_{0}\left(\Phi\left(\tfrac{\Delta/2-l_{\pi_{0}}}{\sigma}\right)-\Phi\left(\tfrac{\Delta/2+l_{\pi_{0}}}{\sigma}\right)\right)\!\!\!\!\!\!\!\!&\!\!\!\!\!\hskip-2.84544pt,\frac{1}{1+\tau}\geq\pi_{0}>0.5\\ (1-\pi_{0})\left(\Phi\left(\tfrac{\Delta/2-l_{\pi_{0}}}{\sigma}\right)-\Phi\left(\tfrac{\Delta/2+l_{\pi_{0}}}{\sigma}\right)\right)\\ +\pi_{0}\left(\Phi\left(\tfrac{l_{\pi_{0}}}{\sigma}\right)-\Phi\left(\tfrac{\Delta-l_{\pi_{0}}}{\sigma}\right)\right)\!\!\!\!\!\!\!\!&\!\!\!\!\!\hskip-2.84544pt,\frac{\tau}{1+\tau}\leq\pi_{0}<0.5\\ \hfil\pi_{0}\!\!\!\!\!\!\!\!&\!\!\!\!\!\hskip-2.84544pt,\pi_{0}<\frac{\tau}{1+\tau}\end{cases}

where τ=fWΔ​(Δ/2)fWΔ​(0)\tau=\tfrac{f_{W}^{\Delta}(\Delta/2)}{f_{W}^{\Delta}(0)} and 𝒞={c1,c2}={h0−lπ0,h0+lπ0}\mathcal{C}=\{c_{1},c_{2}\}=\{h_{0}-l_{\pi_{0}},h_{0}+l_{\pi_{0}}\}. Note that the probability of error with the MAP decision rule depends on Δ/σ\Delta/\sigma and also on π0\pi_{0} (Fig. 2). For smaller values of Δ/σ\Delta/\sigma, the rule yields higher probability of error, which is caused by the information that is lost by the modulo wrapping operation.

IV Modulation and Decoding without Prior Information

IV-A ML Decision Rule

If the prior π0\pi_{0} is not available at the receiver, then the ML decision rule δ\delta is

δ⁡(y~n)\displaystyle\delta(\tilde{y}_{n}) ={0,fWΔ​(y~n−h0)≥fWΔ​(y~n−h1)1,fWΔ​(y~n−h0)<fWΔ​(y~n−h1)\displaystyle=\begin{cases}0,\>\>f_{W}^{\Delta}(\tilde{y}_{n}-h_{0})\geq f_{W}^{\Delta}(\tilde{y}_{n}-h_{1})\\ 1,\>\>f_{W}^{\Delta}(\tilde{y}_{n}-h_{0})<f_{W}^{\Delta}(\tilde{y}_{n}-h_{1})\end{cases} (9)

and the resulting probability of error is

Pe​(δ)\displaystyle P_{e}(\delta)\! ={Φ⁡(h1−h02​σ)−Φ⁡(h1−h02​σ+Δ2​σ),h0+h12​σ≥0Φ⁡(h1−h02​σ−Δ2​σ)−Φ⁡(h1−h02​σ),h0+h12​σ<0\displaystyle=\!\begin{cases}\Phi\left(\frac{h_{1}-h_{0}}{2\sigma}\right)-\Phi\left(\frac{h_{1}-h_{0}}{2\sigma}+\frac{\Delta}{2\sigma}\right),&\frac{h_{0}+h_{1}}{2\sigma}\geq 0\\ \Phi\left(\frac{h_{1}-h_{0}}{2\sigma}-\frac{\Delta}{2\sigma}\right)-\Phi\left(\frac{h_{1}-h_{0}}{2\sigma}\right),&\frac{h_{0}+h_{1}}{2\sigma}<0\end{cases} (10)
Fig. 3: Likelihood functions and ML decision regions. Probability of error with ML decision rule, Pe​(δ)P_{e}(\delta), corresponds to the shaded area under the curves

It is evident from Theorem III.1 that h0=−Δ4h_{0}=-\tfrac{\Delta}{4} and h1=Δ4h_{1}=\tfrac{\Delta}{4} minimizes (10), yielding

Pe​(δ)=Φ⁡(Δ4​σ)−Φ⁡(3​Δ4​σ)P_{e}(\delta)=\Phi\left(\frac{\Delta}{4\sigma}\right)-\Phi\left(\frac{3\Delta}{4\sigma}\right) (11)

The decision regions of ML rule are demonstrated in Fig. 3 on the same setup as in Fig. 1. Note that Pe​(δπ0)≤Pe​(δ)P_{e}(\delta_{\pi_{0}})\leq P_{e}(\delta), and the equality holds when π0=0.5\pi_{0}=0.5.

IV-B Prior Estimation

As it is also evident from (11) and from the shaded regions in Fig. 3, blind ML decision rule results in higher probability of error compared to the MAP decision rule. One possible way to decrease the error rate when priors are unknown at the receiver is to use a two-step decision process: First, apply the ML rule and estimate the prior from the decisions. Then, use the MAP decision rule with the estimated prior.

First, note that the decision rule δ⁡(y~i)\delta(\tilde{y}_{i}) yields to a Bernoulli random variable with

P⁡(δ⁡(y~i)=1)=\displaystyle P(\delta(\tilde{y}_{i})=1)= P⁡(δ⁡(y~i)=1|ai=0)​P​(ai=0)\displaystyle P(\delta(\tilde{y}_{i})=1|a_{i}=0)P(a_{i}=0)
+P⁡(δ⁡(y~i)=1|ai=1)​P​(ai=1)\displaystyle+P(\delta(\tilde{y}_{i})=1|a_{i}=1)P(a_{i}=1)
=\displaystyle= π0​Pe​(δ)+(1−π0)​(1−Pe​(δ)≔pbCLOSE\displaystyle\pi_{0}P_{e}(\delta)+(1-\pi_{0})(1-P_{e}(\delta)\coloneqq p_{b}

and the sum of ML decisions is a binomial random variable, denoted as Nδ=∑i=1Nδ⁡(y~i)∼Binom​(pb)N_{\delta}=\sum_{i=1}^{N}\delta(\tilde{y}_{i})\sim\text{Binom}(p_{b}). Note that

𝔼⁡[Nδ]=N​pb=N⁡(π0​Pe​(δ)+(1−π0)​(1−Pe​(δ))CLOSE\mathbb{E}[N_{\delta}]=Np_{b}=N\left(\pi_{0}P_{e}(\delta)+(1-\pi_{0})(1-P_{e}(\delta)\right)

and the prior π0\pi_{0} can be estimated empirically as

π^0\displaystyle\hat{\pi}_{0} =1N​∑i=1Nδ⁡(y~i)−(1−Pe​(δ))2​Pe​(δ)−1,\displaystyle=\dfrac{\tfrac{1}{N}\sum_{i=1}^{N}\delta(\tilde{y}_{i})-(1-P_{e}(\delta))}{2P_{e}(\delta)-1},

which results in the decision rule

δπ^0​(y~n)\displaystyle\delta_{\hat{\pi}_{0}}(\tilde{y}_{n}) ={0,π^0​fWΔ​(y~n−h0)≥(−π^0)​fWΔ​(y~n−h1)1,π^0​fWΔ​(y~n−h0)<(−π^0)​fWΔ​(y~n−h1)\displaystyle\!=\!\begin{cases}0,\hat{\pi}_{0}f_{W}^{\Delta}(\tilde{y}_{n}\!-\!h_{0})\!\geq\!(1\!-\!\hat{\pi}_{0})f_{W}^{\Delta}(\tilde{y}_{n}\!-\!h_{1})\\ 1,\hat{\pi}_{0}f_{W}^{\Delta}(\tilde{y}_{n}\!-\!h_{0})\!<\!(1\!-\!\hat{\pi}_{0})f_{W}^{\Delta}(\tilde{y}_{n}\!-\!h_{1})\end{cases} (12)

and yields Pe​(δπ^0)P_{e}(\delta_{\hat{\pi}_{0}}).

Fig. 4: Real (π0\pi_{0}) and estimated (π^0\hat{\pi}_{0}) priors using N={200,1000}N=\{200,1000\} symbols (Δ/σ=5\Delta/\sigma=5)

V Experiments

A set of simulated experiment are performed in order to demonstrate the decoding performance of different decision rules. In each experiment, NN random data bits are generated from a Bernoulli distribution Ber​(π0)\text{Ber}(\pi_{0}) and mapped to symbols according to (1). Then, AWGN samples with 𝒩⁡(0,1)\mathcal{N}(0,1) are added to each symbol. Because Δ/σ\Delta/\sigma ratio determines the performance, σ\sigma is fixed and modulo wrapping is performed with different Δ\Delta values. For decoding the noisy, modulo-wrapped symbols, MAP, ML and estimated decision rules in (3), (9) and (12) are used, respectively. The experiments are repeated M=50M=50 times, unless otherwise noted.

In Fig. 4, the estimated priors from different number of symbols are demonstrated. The lines correspond to the average and the shaded regions correspond to the standard deviation of error rate for MM experiments. The plots are limited to π0∈[0,0.5]\pi_{0}\in\left[0,0.5\right] because of symmetry.

Fig. 5: Empirical error rate with different decision rules (Δ/σ=5\Delta/\sigma=5)

In Fig. 5, the resulting error rate of three decision rules for Δ/σ=5\Delta/\sigma=5 are compared. The error rate of the estimated decision rule is close to MAP rule, which is in line with the theoretical probability of error in (8). Both MAP rules with known and estimated priors yield lower error rates than ML rule when the bits are further from uniform, and they yield the same result as expected when π0=0.5\pi_{0}=0.5.

Although the MAP rule with estimated priors results in improved probability of error compared to ML rule for Δ/σ=5\Delta/\sigma=5, it is important to note that its performance inherently depends on the ML rule. If the decisions from ML rule is unreliable, e.g. Δ/σ\Delta/\sigma is too small, then the prior estimate and hence the estimated MAP rule will not perform as well. However, the ratio Δ/σ\Delta/\sigma being too small results in high decoding errors even when the priors are known, as it is seen in Fig. 2. Therefore, operating in low Δ/σ\Delta/\sigma regime is not ideal to begin with.

VI Conclusion

Modulo receivers can provide several improvements in a communication problem. In this work, the optimum communication scheme through a modΔ\md AWGN channel is explored under different scenarios for the available information at the receiver. It is demonstrated that the optimum binary signaling on a modΔ\md AWGN channel to minimize the probability of decoding error is to use 𝒮={−Δ2,Δ2}\mathcal{S}=\{-\tfrac{\Delta}{2},\tfrac{\Delta}{2}\} regardless of the information available at the receiver. This is in line with nested lattice codes in [6], where Λ=Δ​ℤ\Lambda=\Delta\mathbb{Z} and Λ′=Δ2​ℤ\Lambda^{\prime}=\tfrac{\Delta}{2}\mathbb{Z}.

Moreover, a modified, iterative decision rule is proposed instead of ML rule when the prior information is not available at the receiver. Although the modified decision rule might not achieve uniformly better performance than blind estimation, it improves performance when modΔ\md channel is favorable.

References

  • [1] M. Tomlinson, “New automatic equaliser employing modulo arithmetic,” Electronics letters, vol. 7, no. 5, pp. 138–139, 1971.
  • [2] U. Erez, R. Zamir, and S. Member, “Achieving 0.5log(1+SNR) on the AWGN Channel With Lattice Encoding and Decoding,” October, vol. 50, no. 10, pp. 2293–2314, 2004.
  • [3] C. Ling and J. C. Belfiore, “Achieving AWGN channel capacity with lattice Gaussian coding,” IEEE Transactions on Information Theory, vol. 60, no. 10, pp. 5918–5929, 2014.
  • [4] O. Ordentlich, G. Tabak, P. K. Hanumolu, A. C. Singer, and G. W. Wornell, “A modulo-based architecture for analog-to-digital conversion,” IEEE Journal on Selected Topics in Signal Processing, 2018.
  • [5] G. David Forney, M. D. Trott, and S. Y. Chung, “Sphere-bound-achieving coset codes and multilevel coset codes,” IEEE Transactions on Information Theory, 2000.
  • [6] R. Zamir, “Lattices are everywhere,” Information Theory and Applications Workshop, ITA 2009, pp. 392–421, 2009.

Appendix A Proof of Lemma 1

Proof.

By definition, z=fWΔ​(c)=∑k=−∞∞fW​(c+k​Δ)=∑k=−∞∞fW​(−c−k​Δ)=∑k=∞−∞fW​(−c+k​Δ)=∑k=−∞∞fW​(−c+k​Δ)=fΔ​(−c)z=f_{W}^{\Delta}(c)=\sum_{k=-\infty}^{\infty}f_{W}(c+k\Delta)=\sum_{k=-\infty}^{\infty}f_{W}(-c-k\Delta)=\sum_{k=\infty}^{-\infty}f_{W}(-c+k\Delta)=\sum_{k=-\infty}^{\infty}f_{W}(-c+k\Delta)=f^{\Delta}(-c), meaning −c∈ℋ-c\in\mathcal{H}. If 0∉ℋ0\notin\mathcal{H}, there has to be even number of solutions and if there is odd number of solutions, then 0∈ℋ0\in\mathcal{H}. Now assume there are more than two solutions. Without loss of generality, one can assume 0≤c0<c1<Δ/20\leq c_{0}<c_{1}<\Delta/2. Then, by Rolle’s theorem, ∃y∈[c0,c1]:(fΔ)′​(y)=0\exists y\in[c_{0},c_{1}]\colon\left(f^{\Delta}\right)^{\prime}(y)=0. Notice that the derivative (fΔ)′​(y)=∑k=−∞∞fW′​(y+k​Δ)\left(f^{\Delta}\right)^{\prime}(y)=\sum_{k=-\infty}^{\infty}f^{\prime}_{W}(y+k\Delta) can be written as a sum over m={0,…,∞}m=\{0,\dots,\infty\} by pairing sum terms of {(k=m,k=−m−1)}\{(k=m,k=-m-1)\}. Then, (y+k​Δ)​e−1σ2​(y+k​Δ)2+(y−(k+1)​Δ)​e−1σ2​(y−(k+1)​Δ)2<(2​y−Δ)​e−1σ2​(y+k​Δ)2<0\left(y+k\Delta\right)e^{-\tfrac{1}{\sigma^{2}}(y+k\Delta)^{2}}+\left(y-(k+1)\Delta\right)e^{-\tfrac{1}{\sigma^{2}}(y-(k+1)\Delta)^{2}}<\left(2y-\Delta\right)e^{-\tfrac{1}{\sigma^{2}}(y+k\Delta)^{2}}<0, which means (fΔ)′​(y)>0\left(f^{\Delta}\right)^{\prime}(y)>0, which contradicts with Rolle’s theorem. Then, there cannot be more than two solutions. ∎

Appendix B Proof of Lemma 2

Proof.

If aa is uniform over ℤ2\mathbb{Z}_{2} then, L⁡(a=0|c)=L⁡(a=1|c)⇒P⁡(c|a=0)=P⁡(c|a=1)L(a=0|c)=L(a=1|c)\Rightarrow P(c|a=0)=P(c|a=1). Using Poisson’s summation formula,

P⁡(c|a=0)\displaystyle P(c|a=0) =2Δ​∑k=−∞∞e−2​π2​σ2​k2Δ2​cos⁡(2​π​k​c−h0Δ)\displaystyle=\dfrac{2}{\Delta}\sum_{k=-\infty}^{\infty}e^{-\tfrac{2\pi^{2}\sigma^{2}k^{2}}{\Delta^{2}}}\cos\left(2\pi k\frac{c-h_{0}}{\Delta}\right)
P⁡(c|a=1)\displaystyle P(c|a=1) =2Δ​∑k=−∞∞e−2​π2​σ2​k2Δ2​cos⁡(2​π​k​c−h1Δ)\displaystyle=\dfrac{2}{\Delta}\sum_{k=-\infty}^{\infty}e^{-\tfrac{2\pi^{2}\sigma^{2}k^{2}}{\Delta^{2}}}\cos\left(2\pi k\frac{c-h_{1}}{\Delta}\right)

which are equal when c=n2​k​Δ+h0+h12,n,k∈ℤc=\frac{n}{2k}\Delta+\frac{h_{0}+h_{1}}{2},\>n,k\in\mathbb{Z}. Note that f0Δ​(x)=fΔ​(x−h0)f_{0}^{\Delta}(x)=f^{\Delta}(x-h_{0}) is Δ−\Delta-periodic and symmetric around h0h_{0}. Then, f0Δ​((2​h0−c)​modΔ)=f0Δ​((c)​modΔ)f_{0}^{\Delta}(\left(2h_{0}-c\right)\md)=f_{0}^{\Delta}(\left(c\right)\md) == f0Δ​((c+h0−h1)​modΔ)f_{0}^{\Delta}(\left(c+h_{0}-h_{1}\right)\md) == f0Δ​((h0+h1−c)​modΔ)=ρf_{0}^{\Delta}(\left(h_{0}+h_{1}-c\right)\md)=\rho. Lemma 1 entails two of the inputs of these four equations to be equal, which, combined with the constraint on cc, yields h0+h12\tfrac{h_{0}+h_{1}}{2} and h0+h12+sgn⁡(h0+h12)​Δ2\tfrac{h_{0}+h_{1}}{2}+\sgn\left({\frac{h_{0}+h_{1}}{2}}\right)\tfrac{\Delta}{2}. ∎

Appendix C Proof of Theorem III.1

Proof.

Since ana_{n} are independent, the probability of sequence error Pe=1−∏P⁡(a^n=an)P_{e}=1-\prod P(\hat{a}_{n}=a_{n}) is minimized when probability of bit error P⁡(a^n≠an)P(\hat{a}_{n}\neq a_{n}) is minimized for all nn. More formally,

minh0,h1⁡Pe=1−∏(1−minh0,h1⁡P⁡(a^n≠an))\min_{h_{0},h_{1}}P_{e}=1-\prod\left(1-\min_{h_{0},h_{1}}P(\hat{a}_{n}\neq a_{n})\right)

Without loss of generality, assume h0<h1h_{0}<h_{1} and define ℛ0={y~∈[−Δ2,Δ2)|L⁡(a|y~)≥1}\mathcal{R}_{0}=\{\tilde{y}\in\itv\>\big|\>L(a|\tilde{y})\geq 1\} and ℛ1={y~∈[−Δ2,Δ2)|L⁡(a|y~)<1}\mathcal{R}_{1}=\{\tilde{y}\in\itv\>\big|\>L(a|\tilde{y})<1\}. From Lemma 2, it can be seen that

  • •

    If h0+h12<0\tfrac{h_{0}+h_{1}}{2}<0, then

    c0=h0+h12,c1=h0+h12+Δ2\displaystyle c_{0}=\tfrac{h_{0}+h_{1}}{2},c_{1}=\tfrac{h_{0}+h_{1}}{2}+\tfrac{\Delta}{2}
    ℛ1=(c0,c1),ℛ0=[−Δ2,c0]∪[c1,Δ2)\displaystyle\mathcal{R}_{1}=(c_{0},c_{1}),\mathcal{R}_{0}=\left[-\tfrac{\Delta}{2},c_{0}\right]\cup\left[c_{1},\tfrac{\Delta}{2}\right)
  • •

    If h0+h12≥0\tfrac{h_{0}+h_{1}}{2}\geq 0, then

    c0=h0+h12−Δ2,c1=h0+h12\displaystyle c_{0}=\tfrac{h_{0}+h_{1}}{2}-\tfrac{\Delta}{2},c_{1}=\tfrac{h_{0}+h_{1}}{2}
    ℛ0=(c0,c1),ℛ1=[−Δ2,c0]∪[c1,Δ2)\displaystyle\mathcal{R}_{0}=(c_{0},c_{1}),\mathcal{R}_{1}=\left[-\tfrac{\Delta}{2},c_{0}\right]\cup\left[c_{1},\tfrac{\Delta}{2}\right)

Then, the probability of bit error becomes

Pe=\displaystyle P_{e}= P⁡(a^n=1|an=0)​P​(an=0)\displaystyle P(\hat{{a}}_{n}=1|a_{n}=0)P(a_{n}=0)
+P⁡(a^n=0|an=1)​P​(an=1)\displaystyle+P(\hat{{a}}_{n}=0|a_{n}=1)P(a_{n}=1)
=\displaystyle= ∑k=−∞∞(Q⁡(η+k​Δσ)−Q⁡(η+Δ2+k​Δσ))\displaystyle\sum_{k=-\infty}^{\infty}\left(Q\left(\frac{\eta+k\Delta}{\sigma}\right)-Q\left(\frac{\eta+\frac{\Delta}{2}+k\Delta}{\sigma}\right)\right)

where η=h1−h02\eta=\tfrac{h_{1}-h_{0}}{2}. Setting ∂Pe∂η=0\frac{\partial P_{e}}{\partial\eta}=0 yields

∑k=−∞∞fW​(η+Δ2+k​Δ)=∑k=−∞∞fW​(η+k​Δ)\sum_{k=-\infty}^{\infty}f_{W}\left(\eta+\frac{\Delta}{2}+k\Delta\right)=\sum_{k=-\infty}^{\infty}f_{W}\left(\eta+k\Delta\right) (13)

Lemma 2 dictates that the solutions to (13) are in the form of 2​η+Δ2=k​Δ2,k∈ℤ2\eta+\tfrac{\Delta}{2}=k\tfrac{\Delta}{2},k\in\mathbb{Z}, and only 2​η=h1−h0=Δ22\eta=h_{1}-h_{0}=\tfrac{\Delta}{2} satisfies the conditions for h0h_{0} and h1h_{1}. For such h0h_{0} and h1=h0+Δ2h_{1}=h_{0}+\tfrac{\Delta}{2}, the average power becomes P=𝔼⁡[xn2]=12​h02+12​(h0+Δ2)2P=\mathbb{E}[x_{n}^{2}]=\tfrac{1}{2}h_{0}^{2}+\tfrac{1}{2}(h_{0}+\tfrac{\Delta}{2})^{2}. Then, h0∗h_{0}^{*} and h1∗h_{1}^{*} that minimize the average power is h0∗=−Δ4​ and ​h1∗=Δ4h_{0}^{*}=-\frac{\Delta}{4}\text{ and }h_{1}^{*}=\frac{\Delta}{4}. ∎

Appendix D Proof of Theorem III.2

Proof.

Without loss of generality, assume h1=h0+ϵh_{1}=h_{0}+\epsilon, 0≤ϵ<Δ2−h00\leq\epsilon<\tfrac{\Delta}{2}-h_{0} and c1≤c2c_{1}\leq c_{2}. Then, one can see that argmaxϵPe=0\argmax_{\epsilon}P_{e}=0 and argminϵPe=Δ2\argmin_{\epsilon}P_{e}=\frac{\Delta}{2}, by setting

∂Pe∂ϵ=±π0​(fWΔ​(c2−h0)−fWΔ​(c1−h0))=0\dfrac{\partial P_{e}}{\partial\epsilon}=\pm\pi_{0}\left(f_{W}^{\Delta}(c_{2}-h_{0})-f_{W}^{\Delta}(c_{1}-h_{0})\right)=0

and by incorporating the following definition of c1c_{1} and c2c_{2},

fWΔ​(c2−h0)\displaystyle f_{W}^{\Delta}(c_{2}-h_{0}) =fWΔ​(c1−h0)\displaystyle=f_{W}^{\Delta}(c_{1}-h_{0})
fWΔ​(c2−h1)\displaystyle f_{W}^{\Delta}(c_{2}-h_{1}) =fWΔ​(c1−h1)\displaystyle=f_{W}^{\Delta}(c_{1}-h_{1})

into the derivative. Then, the h0h_{0} and h1h_{1} that minimize the power is −Δ4-\tfrac{\Delta}{4} and Δ4\tfrac{\Delta}{4}, as in Theorem III.1. ∎